OpenStudy (anonymous):

A system of equations is given below. 2x + 7y = 1 -3x – 4y = 5

OpenStudy (anonymous):

A). multiply the first equation by 3. + 6x +21y = 3 Add to the second equation. in order to do this, one of your variables needs to have the same number so they can cancel. multiply the second equation by 2. -6x -8y= 10 A./B. Now you can add them and solve. 6x + 21y = 3 + -6x - 8y =10 13y = 13 y=1 Now that you know the value of 1 variable, you can solve for the other variable. Just substitute the y=1 in for 1 in the first equation. 6x +21(1) = 3 6x = -18 x = -3 C. Prove that it equals the first equation. Just plug in the values into the original equation. 2x + 7y =1 x = -3 and y = 1 2(-3) +7(1) = 1 1 = 1

OpenStudy (anonymous):

Thank you SO much!! This makes sense!!

OpenStudy (anonymous):

You're welcome! :) I'm glad!